OpenStudy (anonymous):

second derivative of y=ln(x2+1)

OpenStudy (yttrium):

Just derive the equation twice.

OpenStudy (yttrium):

\[y' = \frac{ 2x }{ x^2+1 }\] --> 1st derivative \[y'' = \frac{2(x^2+1) - 2x(2x) }{ (x^2+1)^2 }\] Hence \[y'' = \frac{2 (1-x^2) }{ (x^2+1)^2 }\]